A 0.68-m aluminum bar is held with its length parallel to the east-west direction and dropped from a bridge. Just before the bar hits the river below, its speed is 16 m/s, and the emf induced across its length is 5.7 x 10-4 V. Assuming the horizontal component of the Earth's magnetic field at the location of the bar points directly north, (a) determine the magnitude of the horizontal component of the Earth's magnetic field, and (b) state whether the east end or the west end of the bar is positive.
Answer #1

Iind

(a)

Induced emf = E = BLv

B = E/Lv = 5.7* 10^(-4) / (0.68* 16) = 5.238 * 10^(-5) T

(b)

East end of the bar should be positive in accordance with the Lenz's law.As the rod goes down, the flux associated with it (effective area as shown) increases so that the induced current opposes the earth's magnetic field.